OpenStudy (anonymous):

WHY or HOW , in R3, is (a,b,c) where b=a+c, NOT A VECTOR SUB- SPACE? WHEREAS in R3, (a,b,c) where c=a-b IS A VECTOR SUB SPACE?

OpenStudy (thomas5267):

@ganeshie8 I am currently studying linear algebra but I have no idea what is happening here.

OpenStudy (anonymous):

YEAH, man, I am trying to understand also, because the first one, according to me , is closed under addition, i tested with some examples such as v1= ( 2,5,3) + v2= (5,9,4) to where v1 and v2 are vectors in V a subset of W, of the form (a,b,c) , so v1 + v2 = (7,14,7) which still suit the given operation b=a+c, same with 3(2,5,3) where 3 is scalar k, which gives me (4,10,6) .

OpenStudy (thomas5267):

If it is closed under addition and it is not a linear subspace, then it must be open under scalar multiplication. I don't see how it is opened though.

OpenStudy (anonymous):

i don't really know, honestly

OpenStudy (thomas5267):

Are you sure that the b=a+c set is not a linear subspace? Who told you that it isn't a linear subspace?

OpenStudy (anonymous):

its in the book, ELEMENTARY LINEAR ALGEBRA, Howard Anton ex 4.2 number 1c

OpenStudy (thomas5267):

I suspect something is wrong with the book.

OpenStudy (anonymous):

i dont eve understand it either

OpenStudy (anonymous):

@MERTICH I suspect it's a typo as well - looks like a subspace to me. A note on your work so far: showing that the rule holds for a few examples doesn't establish that the vector is a subspace (unless you find a counter-example, then you only need one to show it's not). Work with the general form. Take two vectors, \((a_1,b_1,c_1)\) and \((a_2,b_2,c_2)\), then adding them yields \((a_1+a_2,b_1+b_2,c_1+c_2)\), which according to the rule is equivalent to \((a_1+a_2,(a_1+a_2)+(c_1+c_2),c_1+c_2)\). Hence the set of vectors is closed under addition. Now take one vector, \((a_1,b_1,c_1)\), and multiply by a scalar \(k\). You get \((ka_1,kb_1,kc_1)=(ka_1,k(a_1+c_1),kc_1)\). Hence the set is also closed under scalar multiplication. Another condition you must satisfy is that the set must contain the zero vector, which occurs for \(a=c=0\). You could claim that the set isn't a subspace if \(a,b,c\) must be non-zero or distinct, but if that were the case then the second set wouldn't be a subspace.

OpenStudy (anonymous):

thank you @sithsandgiggles, and @thomas5627. You have helped me to confirm my sentiments as well